Find 4 consecutive odd intergers whose sum is 216
Answer by ptaylor(1658)

Let x=first
x+2=the second
x+4=the third
And x+6=fourth
Now we are told that these consecutive odd integers add up to 216. So:
x+x+2+x+4+x+6=216 start collecting like terms
4x+12=216 subtract 12 from both sides
4x+12-12=216-12 and this equals:
4x=204 divide both sides by 4
x=51----first integer
x+2=51+2=53 second integer
x+4=51+4=55 third integer
x+6=61+6=57 fourth integer
CK
51+53+55+57=216
216=216

Hope this helps----ptaylor